### v1

@@ 3 Usage Characteristics
In this section, we present our results on how participants used AR laptops during the two-week diary period. Several factors influenced their device usage, including task, physical environment, social interactions, physical comfort, and aesthetics. Below, we discuss these factors and their effects on usage behaviors. In addition, we report on hybrid use cases in which the AR laptop was used simultaneously with physical devices and tasks. We present the frequency of behaviors and remarks based on our closed coding results (Section 2.4) to reflect the importance of a situation rather than an accurate measurement of how often a situation occurs [66]. We use $n$ to denote the number of sessions during which a code was observed, out of 143 total sessions across all participants (e.g., participants collectively watched videos in $n$ = 59 sessions). We use $N$ to denote the number of participants associated with each code (out of 14; e.g., $N$ = 10 participants watched videos at some point throughout the study). Figure 2 summarizes the main themes that we report on in our results.
@@ 3.1 Task
In 39 sessions, participants configured their AR workspace to consist of only a single window (Figure 5, left). This layout was most frequently used for watching video ($n$ = 14, $N$ = 7), browsing websites ($n$ = 10, $N$ = 5), and reading and writing documents ($n$ = 7, $N$ = 3). For productive use cases, several participants ($N$ = 5) reported that this layout helped them focus on a single task. As P5 commented, “I did not feel the need to open more windows…. I would often open multiple windows when I am multitasking.” P11 used a single window layout to read because it was “free from distractions.” For entertainment purposes, like watching a video, participants appreciated the immersiveness of this arrangement ($N$ = 4). As P10 remarked, “I placed [a single] window in the middle to have a sense of a big TV or cinema.”
@@
Participants mainly arranged windows in a way that separated them in space ($n$ = 59, $N$ = 11; Figure 6, left), ensuring that their contents were fully visible. For P8, separating windows facilitated their tasks by allowing them to “see them all at a glance, without having them stacked on top of each other like a conventional laptop does.” Participants also overlapped ($n$ = 37, $N$ = 11; Figure 6, middle) or stacked ($n$ = 26, $N$ = 11; Figure 6, right) windows. Two participants reported that stacking windows facilitated efficient task switching by allowing them to quickly “swap back and forth” between them (P3).
@@ 3.6 Hybrid Usage
In 98 sessions, participants ($N$ = 14) reported using the AR laptop in combination with physical displays (Figure 10), including phones ($n$ = 83, $N$ = 14), computers ($n$ = 28, $N$ = 9), tablets ($n$ = 1, $N$ = 1), watches ($n$ = 4, $N$ = 2), and TVs ($n$ = 2, $N$ = 2).
@@
When using the AR laptop alongside a conventional laptop or desktop computer, some users reported arranging their virtual windows as peripheral displays around a physical laptop ($N$ = 5). For example, while P2 used their physical laptop for a design task, they displayed reference materials, such as a list of design requirements, as windows in AR and arranged them around their device. P4 used a similar workflow, organizing reference materials around their computer monitor, which they used to author emails. Two participants used the AR laptop alongside physical devices due to the hardware limitations of the former, including its field of view and resolution. They also relied on physical laptops to access applications and information that was unsupported or missing on the AR laptop ($N$ = 5). Several examples of these applications include 3D graphics software (P1) and code editors (P13).
@@
Figure 10: Hybrid usage — Participants often used the AR laptop in combination with other devices, such as a laptop.
@@ 4 Values behind AR Laptop Usage
Six participants appreciated the in situ nature of their windows. In particular, they believed that it allowed them to maintain a higher level of situational awareness of their physical surroundings ($N$ = 4). For example, P11 regarded the ability to remain situated in and aware of the surroundings as critical for their role as an administrator: “I would absolutely need it. If someone approached me, I would have to greet them.” P12 noted that in a public setting, “it is important to have a sense of what is happening in the surroundings,” which is supported by the device. As discussed in Sections 3.2 and 3.6, the AR laptop also offered opportunities to supplement the user’s task and physical environment with contextually relevant digital information.
@@ 5 Current Challenges with AR Laptop Usage
Interaction limitations
Hardware and software limitations aside, our results revealed several fundamental interaction limitations. While the expanded virtual display space provides benefits, such as improving multi-screen workflows, it also presents challenges for navigation across the display ($N$ = 3). As P8 noted, “the trackpad is too small for such a wide canvas.” P2 similarly commented that they “want to be able to make use of that entire space” but disliked “having to manage that using the mouse interface.”
@@
Lastly, it proved difficult for the device to effectively balance the visibility of the physical and virtual environments. For instance, while several participants ($N$ = 6) found value in the immersiveness of the AR laptop, others reported feeling “overwhelmed” when too many windows were opened (P10, $N$ = 6). Similarly, while participants appreciated the in-situ nature of the displays ($N$ = 6), they also found their physical surroundings distracting ($N$ = 7). Combined, these results highlight several core design tensions in the development of AR systems.
@@ 6 Discussion
In this work, we investigated knowledge workers’ usage and perceptions of AR for everyday tasks and activities. Our findings confirm and extend many prior studies on AR. In this section, we relate these works to our results, identify opportunities for future research, and outline the limitations of our study.
@@ 6.1 How Do People Work in AR?
In our study, participants engaged in tasks typical of knowledge workers (e.g., [45]), including reading and writing documents, web browsing, and checking and composing emails. To support these tasks, participants explored various virtual workspaces. Prior research has documented various patterns of window usage in AR, including the use of the central field of view for primary task content [70, 58] and different window configurations for productivity tasks, such as vertical, horizontal, two-by-two, and personal cockpit layouts [63, 33]. In our in-the-wild study, we observed participants using single windows and side-by-side layouts, which supported focused tasks and cross-referencing, respectively. When more than three windows were used, they often represented applications for different tasks. Similar to the findings of Cho et al. [22], we observed participants blending leisure and work applications within their workspace in these settings. Additionally, consistent with Pavanatto et al. [70] and Lischke et al. [58], for these multi-window layouts, participants organized their multi-window layouts around a central task area, while using the periphery for secondary applications. In addition to previously documented behaviors, we observed that participants used their peripheral screen space to maintain persistent AR windows. They sometimes stacked or overlapped their windows to support interactions such as quickly switching back and forth between window contents. They also adjusted window sizes according to their internal content, sometimes to form factors that are not supported by current screen-based computers. Although much of previous work has focused on static layouts (e.g., [19, 12]), our results indicate that virtual workspaces tend to be dynamically adjusted during usage.
@@
Besides configuring their workspace based on task considerations, participants adapted window placements to account for environmental lighting conditions, as noted by Pavanatto et al. [70]; geometric features and visual clutter, as highlighted by Cheng et al. [19]; and the presence of others, as documented by Ng et al. [67] and Medeiros et al. [64]. Interestingly, participants also considered the perceived spaciousness of their environment in their workspace configurations, a factor that, to our knowledge, has not been previously documented.
@@
While many of our participants used the AR laptop as intended, i.e., as a standalone laptop, we observed several participants shifting to more hybrid forms of usage. As Pavanatto et al. [73] previously showed, combining physical and virtual monitors may offer a better balance of familiarity with physical monitors, while also providing extra space from virtual monitors. Our results showed that such a hybrid usage approach emerged in the wild when using an AR laptop, deriving similar benefits. This highlights the potential for supporting integrated multi-device workflows with AR, rather than solely suggesting that AR should completely replace current devices, as previously proposed in Auda et al. [5]. In addition to complementing physical screens, we also observed participants using the AR device in combination with physical tasks. This similarly echoes prior research suggesting that the unique value of AR lies in its seamless blending of virtual and real (e.g., [57]).
Lastly, the emergent values we observed — such as improved physical comfort, portability, multiscreen workflows, display in situ, immersiveness, and privacy — largely support the idea that the benefits of AR previously noted in controlled laboratory settings (e.g.,  [70, 12, 81, 79]) indeed translate to real-world applications. Our work also reflects several long-standing challenges documented in previous research [49], including some arising from hardware and software issues, as well as others more fundamental to AR interactions.
@@ 6.2 Opportunities
Our results also indicate that in addition to tackling hardware and software challenges, several fundamental interaction issues must be addressed to facilitate the effective daily use of AR. One common issue we observed was efficiently navigating the expanded AR display space. We encourage researchers to explore new methods to improve this process, potentially referencing related literature on large displays (e.g., [27]), where similar interaction challenges have also been identified. Another challenge relates to window management. To address this issue, we believe there is value in operationalizing common arrangement patterns we observed as layout templates that users can snap windows to, similar to the application tiling feature found in macOS [4]. However, given the sensitivity of window arrangements to their surroundings, more intelligent adaptation approaches may be necessary, such as the one previously proposed by Cheng et al. [19]. Furthermore, participants differed in their window arrangement behaviors. We believe that future work should further explore how personalized AR workspaces can use contextual factors to assist users in their tasks.
@@
Despite improvements in hardware, software, and interactions, challenges may still persist in promoting adoption and effective usage. In our study, we found that users tended to revert to their usual window usage habits in physical workspaces. In many cases, these usage patterns did not take advantage of the unique affordances of AR. Participants themselves even reported feeling unsure about how to effectively use the new medium. A possible solution to these challenges could involve improving upon the onboarding process. Although we attempted to thoroughly explain the device functionalities during our introductory session, providing more detailed and repeated training might be helpful. To help users overcome their legacy biases regarding window arrangement patterns in particular, future work could consider exploring context-based and situational onboarding methods [17], such as suggesting layouts tailored to their current task. Yet, while we see AR as a means to engage with digital information in new ways, we also recognize the importance of maintaining familiar interfaces for work to facilitate a smoother transition from current laptops and devices. Future devices should leverage familiar aspects while introducing new features, but this is likely easier said than done, requiring careful design of onboarding procedures and interfaces. Perhaps parallels can be drawn to the introduction of mobile user interfaces. Early mobile user interface designs heavily relied on the preceding desktop paradigm [76]. While these designs did not initially suit the target form factor, they nonetheless provided a familiar entry point for users. AR interfaces today may be in a similar state. As previously discussed by Pavanatto et al. [72], drawing on familiar paradigms may ease the transition from today’s personal computers to computing with virtual displays, thereby enabling users to adapt more quickly and effectively to the new medium.
@@ 6.3 Limitations
Our current study design is subject to several limitations. First, while we aimed to study usage behaviors longitudinally, our results currently reflect the participants’ engagement with the device for only 40 minutes a day over a two-week period. Although we still consider this substantial, with an average of seven hours of use per person, extending the usage duration further would undoubtedly improve the ecological validity and generalizability of our results. In particular, while our study design exposed participants to the AR laptop for an extended duration, we were unable to fully capture the behaviors that may emerge if the device completely replaced their current laptops or computers. While we encourage researchers to study AR usage over longer periods, it is important to acknowledge the significant logistical and feasibility challenges involved. An alternative approach to consider is to consult and gather feedback from current power users.
@@
Furthermore, our study was based on a specific hardware and software implementation of AR, which contains known limitations and is not wholly representative of future AR devices. It is not unlikely that as AR devices improve, user behaviors and perceptions may also change. For example, if the device supported more applications and was better integrated with participants’ individual information ecosystems, they might have used the device more exclusively, rather than in combination with other computers. Replications of our study with future devices would, therefore, yield valuable insights.

### v2

@@ 4 Usage Characteristics
In this section, we present our results on how participants used AR laptops during the two-week diary period. Several factors influenced their device usage, including task, physical environment, social interactions, physical comfort, and aesthetics. In addition, we report on hybrid use cases in which the AR laptop was used simultaneously with physical devices and tasks. We present the frequency of behaviors and remarks based on our closed coding results (Section 3.4) to reflect the importance of a situation rather than an accurate measurement of how often a situation occurs [65]. We use $n$ to denote the number of sessions during which a code was observed, out of 143 total sessions across all participants (e.g., participants collectively watched videos in $n$ = 59 sessions). We use $N$ to denote the number of participants associated with each code (out of 14; e.g., $N$ = 10 participants watched videos at some point throughout the study). Figure 2 summarizes the main themes that we report on in our results.
@@ 4.1 Task
In 39 sessions, participants configured their AR workspace to consist of only a single window (Figure 5, left). This layout was most frequently used for watching video ($n$ = 14, $N$ = 7), browsing websites ($n$ = 10, $N$ = 5), and reading and writing documents ($n$ = 7, $N$ = 3). For productive use cases, several participants ($N$ = 5) reported that this layout helped them focus on a single task. As P5 commented, “I did not feel the need to open more windows…. I would often open multiple windows when I am multitasking.” P11 used a single window layout to read because it was “free from distractions”. For entertainment purposes, like watching a video, participants appreciated the immersiveness of this arrangement ($N$ = 4). As P10 remarked, “I placed [a single] window in the middle to have a sense of a big TV or cinema”.
@@
Participants mainly arranged windows in a way that separated them in space ($n$ = 59, $N$ = 11; Figure 6, left), ensuring that their contents were fully visible. For P8, separating windows facilitated their tasks by allowing them to “see them all at a glance, without having them stacked on top of each other like a conventional laptop does.” Participants also overlapped ($n$ = 37, $N$ = 11; Figure 6, middle) or stacked ($n$ = 26, $N$ = 11; Figure 6, right) windows. P3 and P10 reported that stacking windows facilitated efficient task switching by allowing them to quickly “swap back and forth” between them.
@@ 4.6 Hybrid Usage
Hybrid display
In 98 sessions, participants ($N$ = 14) reported using the AR laptop in parallel with other physical displays (Figure 10), even though the device did not technically support cross-device interactions, as discussed earlier, due to the lack of tracking capabilities and software support for such interactions. In such scenarios, participants manually positioned the virtual windows to fit their needs on the AR laptop, and switched inputs between devices opportunistically. These concurrent uses included phones ($n$ = 83, $N$ = 14), computers ($n$ = 28, $N$ = 9), tablets ($n$ = 1, $N$ = 1), watches ($n$ = 4, $N$ = 2), and TVs ($n$ = 2, $N$ = 2).
@@
When using the AR laptop alongside a conventional laptop or desktop computer, some users reported manually arranging their virtual windows around a physical laptop ($N$ = 5). It is important to note that these arrangements required users to manually switch between the physical device’s and the AR laptop’s input. For example, while P2 used their physical laptop for a design task, they displayed reference materials, such as a list of design requirements, as windows in AR and arranged them around their device. P4 used a similar workflow, organizing reference materials around their computer monitor, which they used to author emails. Two participants used the AR laptop alongside physical devices due to the hardware limitations of the former, including its field of view and resolution. They also relied on physical laptops to access applications and information that was unsupported or missing on the AR laptop ($N$ = 5). Several examples of these applications include 3D graphics software (P1) and code editors (P13).
Hybrid task
@@
Figure 10: Hybrid usage — Participants often used the AR laptop in combination with other devices, such as a laptop. The images shown are created based on participant descriptions, using the same compositing approach as in Augmented Reality Productivity In-the-Wild: A Diary Study of Usage Patterns and Experiences of Working with AR Laptops in Real-World Settingsright.
@@ 5 Values behind AR Laptop Usage
Six participants appreciated the in situ nature of their windows. In particular, they believed that it allowed them to maintain a higher level of situational awareness of their physical surroundings ($N$ = 4). For example, P11 regarded the ability to remain situated in and aware of the surroundings as critical for their role as an administrator: “I would absolutely need it. If someone approached me, I would have to greet them.” P12 noted that in a public setting, “it is important to have a sense of what is happening in the surroundings,” which is supported by the device. As discussed in Sections 4.2 and 4.6, the AR laptop also offered opportunities to supplement the user’s task and physical environment with contextually relevant digital information.
@@ 6.1 Hardware and software
6.2 Interaction limitations
Hardware and software limitations aside, our results revealed several fundamental interaction limitations.
Expanded display space navigation
While the expanded display provides benefits, such as improving multi-screen workflows, it also presents challenges for navigation across the display ($N$ = 3). As P8 noted, “the trackpad is too small for such a wide canvas.” P2 similarly commented that they “want to be able to make use of that entire space” but disliked “having to manage that using the mouse interface.”
Window management
@@ 6.2 Interaction limitations
Balancing virtual and physical
Lastly, it proved difficult for the device to effectively balance the visibility of the physical and virtual environments. For instance, while several participants ($N$ = 6) found value in the immersiveness of the AR laptop, others reported feeling “overwhelmed” when too many windows were opened (P10, $N$ = 6). Similarly, while participants appreciated the in situ nature of the displays ($N$ = 6), they also found their physical surroundings distracting ($N$ = 7). Combined, these results highlight several core design tensions in the development of AR systems.
@@ 7 Discussion
In this work, we investigated knowledge workers’ usage and perceptions of AR for everyday tasks and activities. Our findings confirm and extend many prior studies on AR. In this section, we relate these works to our results, discuss emergent usage patterns, identify opportunities for future research, and outline the limitations of our study.
@@ 7.1 How Do People Work in AR?
In our study, participants engaged in tasks typical of knowledge workers (e.g., [42]), including reading and writing documents, web browsing, and checking and composing emails. To support these tasks, participants explored various virtual workspaces. Prior research has documented various patterns of window usage in AR, including the use of the central field of view for primary task content [70, 55] and different window configurations for productivity tasks, such as vertical, horizontal, two-by-two, and personal cockpit layouts [62, 29]. In our in-the-wild study, we observed participants using single windows and side-by-side layouts, which supported focused tasks and cross-referencing, respectively. When more than three windows were used, they often represented applications for different tasks. Similar to the findings of Cho et al. [21], we observed participants blending leisure and work applications within their workspace in these settings. Additionally, consistent with Pavanatto et al. [70] and Lischke et al. [55], for these multi-window layouts, participants organized their multi-window layouts around a central task area, while using the periphery for secondary applications.
Besides previously documented behaviors, we observed that participants used their peripheral screen space to maintain persistent AR windows. They sometimes stacked or overlapped their windows to support interactions such as quickly switching back and forth between window contents. They also adjusted window sizes according to their internal content, sometimes to form factors that are not supported by current screen-based computers. Although much of previous work has focused on static layouts (e.g., [20, 11]), our results indicate that virtual workspaces tend to be dynamically adjusted during usage.
@@
Besides configuring their workspace based on task considerations, participants adapted window placements to account for environmental lighting conditions, as noted by Pavanatto et al. [70]; geometric features and visual clutter, as highlighted by Cheng et al. [20]; and the presence of others, as documented by Ng et al. [66] and Medeiros et al. [63]. Participants also considered the perceived spaciousness of their environment in their workspace configurations, a factor that has not been previously documented.
@@
One interesting discrepancy we observed regarding the environment is that while some participants valued the device’s support for more situated usage, others reported that the in situ display of virtual content obscured their physical surroundings. We believe this apparent contradiction stems from the inherent trade-off in AR between presenting virtual information and maintaining unobstructed real-world awareness. Presenting information in situ can support a more seamless integration of virtual and physical tasks; however, introducing additional visual information will inevitably block some of a user’s view of their physical environment. Depending on the scenario, this obstruction can be more or less acceptable. For example, if a user is passively working within a physical space, they may find that having virtual windows integrated into their environment offers an appropriate level of situational awareness, allowing them to feel present while attending to virtual tasks. However, if they are instead participating in a collaborative discussion, the visual occlusion can become a more significant impediment to their situational awareness and social interaction.
@@
While many of our participants used the AR laptop as intended, i.e., as a standalone laptop, we observed several participants adopting hybrid forms of usage. As Pavanatto et al. [73] previously showed, combining physical and virtual monitors may offer a better balance of familiarity with physical monitors, while also providing extra space from virtual monitors. Our results showed that such a hybrid usage approach emerged in the wild when using an AR laptop, deriving similar benefits. This highlights the potential for supporting integrated multi-device workflows with AR, rather than solely suggesting that AR should completely replace current devices, as previously proposed in Auda et al. [5]. In addition to complementing physical screens, we also observed participants using the AR device in combination with physical tasks. This similarly echoes prior research suggesting that the unique value of AR lies in its seamless blending of virtual and real (e.g., [54]).
AR versus physical displays
Our study highlights several advantages of AR over physical displays in the wild, including improved physical comfort, portability, multiscreen workflows, in situ display, immersiveness, and privacy. These advantages are largely enabled by AR’s affordances, which support unique usage patterns. One notable pattern facilitated by AR’s expanded display is the utilization of peripheral spaces to host persistent windows and support multiple parallel tasks, a capability previously limited by the finite screen real estate of physical displays. Similarly, AR’s flexibility in workspace arrangements, a contrast to the fixed nature of physical screens, enabled users to design window layouts for new usage postures, such as reclining in bed. Last but not least, by enabling the in situ display of virtual content, we observed AR supporting hybrid workflows that allowed users to simultaneously perform physical tasks and access virtual information. Unlike physical displays, which demanded either the user’s hands or dedicated space, AR’s less intrusive, hands-free display approach provided more freedom in accomplishing diverse physical activities.
In our study, we also observed notable similarities in usage patterns. For instance, many workspace arrangements mirrored common screen-based window layouts. Similarly, while the device had an expanded display space, our results also indicate that users commonly designed workspaces that contained a small number of windows (i.e., 2.4 windows on average), which could also have been supported by physical display setups. We speculate that this may be due to legacy biases, system limitations, and the nature of the sort of tasks our users performed. First, the AR laptop is largely designed as a display extension rather than a fundamentally different interface. This may have prompted users to approach it with a mental model similar to their existing devices. Second, the AR laptop’s hardware limitations, notably its limited field of view, could have restricted the range of workspace usage. Furthermore, because the AR laptop exclusively supported 2D browser-based applications, users could not fully explore more AR-specific use cases. Finally, for the everyday knowledge-work oriented usage scenarios, it is possible that users did not inherently need to leverage AR’s unique affordances, such as support for a greater number of virtual windows.
Overall, while the AR laptop enables several new and unique workflows, users also adopted usage patterns that sometimes mirrored those established with physical displays.
@@ 7.2 Opportunities
Our results also indicate that in addition to tackling hardware and software challenges, several fundamental interaction issues must be addressed to facilitate the effective daily use of AR. One common issue we observed was efficiently navigating the expanded AR display space. Efforts within the current literature have explored ways to support more efficient navigation, including the use of gaze as an additional input modality [12]. We encourage researchers to build on these efforts, particularly by studying ways to improve this process outside controlled laboratory settings. Another challenge relates to window management. To address this issue, we believe there is value in operationalizing common arrangement patterns we observed as layout templates that users can snap windows to, similar to the application tiling feature found in macOS [4]. However, given the sensitivity of window arrangements to their surroundings, more intelligent adaptation approaches may be necessary. Several prior works have explored optimization-based approaches to adapting interfaces based on various contextual factors, such as environment semantics [20], user placement preferences [67], and usability requirements [30]. Building on Lu and Xu’s findings [60], we speculate that future interfaces will need to support a combination of manual arrangement and automatic adaptation. We encourage future work to continue investigating such mixed-initiative approaches.
@@
Despite improvements in hardware, software, and interactions, challenges may still persist in promoting adoption and effective usage. As discussed in Section 7.1, we found that users tended to revert to their usual window usage habits in physical workspaces. In many cases, these usage patterns did not take advantage of the unique affordances of AR. Participants themselves even reported feeling unsure about how to effectively use the new medium. A possible solution to these challenges could involve improving upon the onboarding process. Although we attempted to thoroughly explain the device functionalities during our introductory session, providing more detailed and repeated training might be helpful. To help users overcome their legacy biases regarding window arrangement patterns in particular, future work could consider exploring context-based and situational onboarding methods [16], such as suggesting layouts tailored to their current task. Yet, while we see AR as a means to engage with digital information in new ways, we also recognize the importance of maintaining familiar interfaces for work to facilitate a smoother transition from current laptops and devices. Future devices should leverage familiar aspects while introducing new features, but this is likely easier said than done, requiring careful design of onboarding procedures and interfaces. Perhaps parallels can be drawn to the introduction of mobile user interfaces. Early mobile user interface designs heavily relied on the preceding desktop paradigm [76]. While these designs did not initially suit the target form factor, they nonetheless provided a familiar entry point for users. AR interfaces today may be in a similar state. As previously discussed by Pavanatto et al. [72], drawing on familiar paradigms may ease the transition from today’s personal computers to computing with virtual displays, thereby enabling users to adapt more quickly and effectively to the new medium.
@@ 7.3 Limitations
Our current study design is subject to several limitations. First, while we aimed to study usage behaviors longitudinally, our results currently reflect the participants’ engagement with the device for only 40 minutes a day over a two-week period. It should also be noted that participants received a monetary incentive to complete their sessions, which may have introduced biases in how and when they chose to use the device. To capture behaviors that may emerge if the device completely replaces users’ current laptops or computers, we encourage researchers to continue studying AR usage over longer periods. However, it is also it is important to acknowledge the significant logistical and feasibility challenges involved. An alternative approach to consider is to consult and gather feedback from current power users.
@@
Furthermore, our study was based on a specific hardware and software implementation of AR, which contains known limitations and is not wholly representative of future AR devices. It is not unlikely that as AR devices improve, user behaviors and perceptions may also change. For example, if the device supported more applications and was better integrated with participants’ individual information ecosystems, they might have used the device more exclusively, rather than in combination with other computers. Similarly, with an expanded field of view, participants may be more inclined toward using the peripheral spaces within their display. With greater adoption, we may also expect social norms to change, encouraging more public usage of the device and a distancing from legacy biases and current aesthetic preferences. However, we believe many of our findings will also persist. For example, we expect that comfort will remain a key consideration in workspace design. Specifically, we speculate that certain window arrangements, such as single-window or side-by-side usage, will probably remain in use, particularly in contrast to top-and-bottom arrangements, because they facilitate more ergonomic usage. Additionally, if users continue to employ AR laptops for tasks similar to those observed in this study, we speculate that their workspace arrangement strategies will also remain consistent. For instance, for the same tasks, users will likely not change the number of windows used or how they are arranged, as the task structure will likely necessitate the same interactions. Replications of our study with future devices would yield valuable insights into how usage patterns may evolve or persist.